OpenStudy (anonymous):

I need help graphing this function. I keep getting wrong and i'm not sure what to do. f(x)= 4-(x+2)^3. PLEASE HELP!!!!

hero (hero):

@adriana_98, The vertex form of the equation of a parabola is f(x) = a(x - h)^2 + k where (h , k) is the vertex right?

OpenStudy (anonymous):

yes you can send an attachment

OpenStudy (anonymous):

yes that's right

hero (hero):

Actually, I just noticed the equation is cubed....

OpenStudy (anonymous):

yeah sorry i just saw that

hero (hero):

Well, that's okay, it's still just a transformation of y = x^3

hero (hero):

If you know how to graph y = x^3 then you can do a few horizontal/vertical shifts and reflections to get y = -(x + 2)^3 + 4

OpenStudy (anonymous):

ok, i can do that and for my working out, i would just say where i did the vertical shift and the horizontal shift and dilation?

OpenStudy (anonymous):

@Hero

hero (hero):

You can simply graph it to show such transformations. I would graph both y =x^3 and y = -(x + 2)^3 + 4 in different colors and demonstrate that I arrived at the transformed function by beginning with the parent function.

OpenStudy (anonymous):

ok, thankyou very much for your help